The alkali metals react with the halogens to form ionic metal halides. What mass of potassium chloride forms when 5.11 L of chlorine gas at 0.943 atm and 286 K reacts with 29.0 g potassium?
Answer #1

Answer:

2 K + Cl2 >> 2 KCl

Moles of Cl2= PV/RT (Ideal gas law)

Given for Cl2 gas P=0.943 atm, V=5.11 L, R=0.082 L.atm.K-1 and T= 286 K.
Moles Cl2 = 5.11*0.943/ 0.0821 x 286 =0.205
Moles KCl = 0.205x 2 =0.410 mol KCl
Mass KCl = 0.410 mol x 74.55 g/mol =30.63 g.